OpenStudy (anonymous):

Find the distance between the given points. C(7, 1) and D(4, -6) Distance =

OpenStudy (mertsj):

Do you know the distance formula?

OpenStudy (anonymous):

 The distance needs to fit in the two spaces under the radical. I don't know how to do it.

OpenStudy (anonymous):

The information I gave for the question is what was given :/

OpenStudy (mertsj):

\[d=\sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{1})^2}\]

OpenStudy (mertsj):

That's the distance formula. It says to put the x values of your ordered pairs in and the y values of your ordered pairs in. Then subtract them. Then square. Can you do that?

OpenStudy (mertsj):

Your x values are 7 and 4

OpenStudy (mertsj):

Your y values are 1 and -6

OpenStudy (anonymous):

Honestly, it's all confusing :/ I am homeschooling and I was never taught this. Ok so I then I would subtract them?

OpenStudy (mertsj):

Well put the numbers in and see what happens. You can post your result and we'll see how it goes.

OpenStudy (anonymous):

Ok, give me one second :)

OpenStudy (anonymous):

Ok, I subtracted the 7 from 4 and multiplied 3X3=9 then I subtracted 1-neg6=-5 multiplied -5X5? and got 25. I did it wrong huh?

OpenStudy (mertsj):

But 1-(-6) is 1+6 which is 7

OpenStudy (anonymous):

ohh! ok so it would be 7X7=49 + the 9 = 58 as the answer?

OpenStudy (mertsj):

So now we have: \[d=\sqrt{9+49}=\sqrt{58}\]

OpenStudy (mertsj):

Don't forget the radical.

OpenStudy (anonymous):

Thank You for the help! I understand it now!

OpenStudy (mertsj):

Do you know the Pythagorean theorem?

OpenStudy (mertsj):

Because if you do, I want to show you something.

OpenStudy (anonymous):

Once I see it, I will remember.

OpenStudy (mertsj):

a^2+b^2=c^2. It's for right triangles.

OpenStudy (anonymous):

Oh yes, I know those

OpenStudy (mertsj):

Ok. Watch this.

OpenStudy (mertsj):